a 10kg mass is suspended from the end of a spring and released. If the spring oscillates in shm with a period of 0.5s Cal the energy given to the 10kg Mass if the amplitude of oxidation is 2cm?


1
Expert's answer
2022-02-20T15:45:54-0500

Given:

"m=10\:\rm kg"

"T=0.5\:\rm s"

"A=0.02\:\rm m"


The energy of a SHM

"E=\frac{m\omega^2A^2}{2}=\frac{4m\pi^2A^2}{2T^2}"

"E=\frac{4*10*\pi^2*0.02^2}{2*0.5^2}=0.32\:\rm J"